OpenStudy (anonymous):

A 1.000 gram sample of an alcohol was burned in oxygen to produce 1.913 g of CO2 and 1.174 g of H20. The alcohol contained only C, H, and O. What is the simplest formula of the alcohol?

OpenStudy (anonymous):

The formula will be C2H6O.

OpenStudy (anonymous):

how will i solve this my friend?

OpenStudy (anonymous):

1.913 g CO2/44 = 0.043 moles of CO2*12 = 0.043 moles of C 1.174 g H2O/18 = 0.065 moles of H2O or 0.13 moles of H (because it contains 2 H) O: 1- 0.516 g C - 0.13 g H = 0.354 g O/16 = 0.022 moles of O So we have 0.043:0.13:0.022 as C:H:O Dividing by the smallest no. 0.022 we get 1.95:5.9:1 or 2:6:1 as C:H:O

OpenStudy (anonymous):

thanks a lot my friend. i understand our topic more. but can I ask more questions?

OpenStudy (anonymous):

your welcome :) and sure

OpenStudy (anonymous):

consider the reaction 2NaCl + H2SO4 --> 2HCl + Na2SO4 Ung tanong If the actual yield of 02 was 3.0g, what was the percent yield?

OpenStudy (anonymous):

are you still there kunoichi? :(